When convex lens of focal length 20cm is placed in a liquid whose optical density is equal to the lens then the focal length of that lens will be __________

Grade:12th pass

1 Answers

Saurabh Koranglekar
3 years ago
The focal length 'f' is given by the formula:

1/f = { ( ( n2 / n1) - n1) (1 / R1 - 1 / R2) }

where R1 and R2 are the radius of curvatures of the lens (es) with the proper sign convention.

And n1 and n2 are the refractive indices of medium 1 and 2 respectively.

Now, when n1= n2,

We have, 1/f = 0 × (1/ R1 - 1/R2 )

Hence, 1/f = 0

So,focal length f becomes infinite.